OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x + 2y = 6. Answer x-intercept: (0, 2), y-intercept: (0, 3) x-intercept: (2, 0), y-intercept: (0, 3) x-intercept: (3, 0), y-intercept: (2, 0) x-intercept: (0, 3), y-intercept: (0, 2)

OpenStudy (amistre64):

given the form: ax+by = c the intercepts can be read as: x = c/a, y = c/b

OpenStudy (amistre64):

also good to note is that xint has a y=0, and a yint has an x=0